# Quantum Teleportation Coexisting with Classical Communications in Optical Fiber

### v3

@@ 3.2 Quantum teleportation coexisting with high-rate classical communications
Fig. 3 shows four-fold coincidence counts as we scan Bob’s polarization analyzer setting around the Bloch sphere for Alice transmitting $|{D}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle+|{V}\rangle)$ and $|{A}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle-|{V}\rangle)$. The coincidences exhibit sinusoidal fringes that coincide with Alice’s prepared qubit. We obtain fringe visibilities of $V_{D}=81.3\pm 5.4\%$ and $V_{A}=74.7\pm 4.7\%$, both values being above the $V=1/3$ classical limit [52].
Next, we perform maximum likelihood quantum state tomography [53] on Bob’s target photon to reconstruct the density matrix $\rho_{B}$ of his quantum state conditioned on a BSM detection. Fig. 4 shows $\rho_{B}$ for Alice transmitting the states $|H\rangle$, $|V\rangle$, $|D\rangle$, and $|A\rangle$. The fidelity of $\rho_{B}$ to Alice’s ideal qubit $|\psi\rangle_{A}$ is determined by calculating $F_{\psi}=\langle\psi_{A}|\rho_{B}|\psi_{A}\rangle$. State transfer in the H/V basis does not require HOM interference due to the projection onto $|H\rangle_{D_{1}}|V\rangle_{D_{2}}$ in the Bell state analyzer [54], where raw fidelity is mainly limited from unity due to multi-photon pair degradation. In this basis, we measure $F_{H}=97.5\pm 1.2\%$ and $F_{V}=95.8\pm 2.5\%$ for $|{\psi}\rangle_{A}=|{H}\rangle$ and $|{V}\rangle$, respectively. However, transferring coherent superpositions of $|H\rangle$ and $|V\rangle$ requires interference and thus exhibit more decoherence due to an imperfect spectro-temporal indistinguishability between Alice and Bob’s photons. For Alice transmitting $|{\psi}\rangle_{A}=|{D}\rangle$ and $|{A}\rangle$, we obtain fidelities of $F_{D}=87.5\pm 3.9\%$ and $F_{A}=85.5\pm 3.7\%$, respectively. Knowing that SpRS noise rates are independent of Alice’s prepared qubit, assuming symmetry of the equatorial states we calculate the average fidelity for an arbitrary qubit via $F_{\rm avg}=(F_{H}+F_{V}+2(F_{D}+F_{A}))/6$ in which we obtain $F_{\rm avg}=89.9\pm 3.1\%$. Furthermore, we measure $F_{\rm avg}=90.8\pm 0.8\%$ without sending the photons over the 30.2-km link (see supp. materials), demonstrating that there is minimal degradation of fidelity from distribution over the fiber nor the inclusion of the 400-Gbps signal. The state transfer fidelity of each qubit is shown in Fig. 4(e). All are well above the $F=2/3$ limit for classical-based methods [52], demonstrating that non-classical teleportation can be achieved alongside high-rate conventional communications.
@@ 4 Discussion and Outlook
Our results also indicate the potential for integrating even more advanced quantum operations. For example, simple modifications to Alice’s photon pair source would generate entangled photon pairs and result in entanglement swapping [55], entangling initially independent photons at Alice and Bob’s nodes after the BSM. Based on this similarity, it is reasonable to conclude that swapping alongside high-rate classical data is achievable. Further, midpoint BSMs are important for applications such as networking with a central quantum processor [34], quantum repeaters [29], amongst others. Although these will require more capabilities beyond those demonstrated here (eg., quantum memories, purification, etc.), our results are promising and could be useful for the ongoing development of this technology.
Although O-band teleportation networks show promise for high-fidelity integration, C-band quantum networking is usually preferred in quantum-dedicated fiber due to the lowest loss in silica fibers. Unfortunately, C-band channels suffer from orders of magnitude higher SpRS generation rates from C-band classical light [4, 17]. CV quantum encoding can help address this issue [25] but is much more fragile to loss compared to discrete-variables which limits its reach. One intriguing possibility is to use high-dimensional entanglement [56], but has yet to be fully investigated in high-power regimes. Using a theoretical model [48], we predict our teleportation fidelity would be significantly impacted with mW-level C-band classical power if we used C-band quantum signals (see supp. materials), but coincidence rates would roughly triple. This has relevance for research towards long-distance quantum backbone networks since conventional backbone fibers can have $\sim 100$ mW aggregate power. Further, WDM bandwidth may become limited as some conventional networks push towards fully populating the C- and L-bands [57], which also increases the difficulty of avoiding channel cross-talk and optical amplifier noise. One route to alleviate lower O-band quantum rates is to use wavelength multiplexing, which has been shown successful in allowing multi-channel entangled photons populating most of the O-band co-propagating with high-power C- and L-band WDM channels [17]. In any case, future research into novel methods to protect quantum fidelity from strong background noise would be invaluable.
@@
It is interesting to consider whether teleporting quantum information has any unique benefits in quantum-classical networking compared to direct transmission. Interestingly, we predict that our system can achieve higher coexisting classical power compared to if Alice had directly sent her qubit to Bob over the fiber using a WCS or a heralded single-photon (see supp. materials). This could be due to the stronger correlations in detection across the four photons in our system, and also the segmenting of the longer fiber with a midpoint BSM [28]. Future studies will compare with other teleportation designs such as those using a WCS source at Alice’s node [33, 36, 39] instead of our heralding approach. Looking beyond state teleportation, it would be interesting to investigate the impact of coexisting classical channels on concatenated entanglement swapping or quantum repeaters.
In regards to real-world deployment, classical light-based signals are often used in quantum networking for synchronizing distant pulsed quantum sources and detectors to a common clock, communicating measurement results (eg., BSM results), and monitoring environmental disturbances. Similarly, multiplexing these signals can enhance functionality without sacrificing potentially limited fiber resources [16, 59, 22, 21]. Since these can typically operate at sub-mW powers, our results show that including these in our system would have negligible impact if allocated to the C- or L-bands.

### v4

@@ 3.2 Quantum teleportation coexisting with high-rate classical communications
Fig. 3 shows four-fold coincidence counts as we scan Bob’s polarization analyzer setting around the Bloch sphere for Alice transmitting $|{D}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle+|{V}\rangle)$ and $|{A}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle-|{V}\rangle)$. We observe for each case a sinusoidal fringe that resembles Alice’s prepared qubit with a maximum four-fold rate of $\sim$0.09 counts per second. We obtain fringe visibilities of $V_{D}=81.3\pm 5.4\%$ and $V_{A}=74.7\pm 4.7\%$, both values being above the $V=1/3$ classical limit [52].
Next, we perform maximum likelihood quantum state tomography [53] on Bob’s target photon to reconstruct the density matrix $\rho_{B}$ of his quantum state conditioned on a BSM detection. Fig. 4 shows $\rho_{B}$ for Alice transmitting the states $|H\rangle$, $|V\rangle$, $|D\rangle$, and $|A\rangle$. The fidelity of $\rho_{B}$ to Alice’s ideal qubit $|\psi\rangle_{A}$ is determined by calculating $F_{\psi}=\langle\psi_{A}|\rho_{B}|\psi_{A}\rangle$. State transfer in the $H/V$ basis does not require HOM interference due to the projection onto the qubit eigen-states in the Bell state analyzer [54], where raw fidelity is mainly limited from unity due to multi-photon pair degradation. In this basis, we measure $F_{H}=97.5\pm 1.2\%$ and $F_{V}=95.8\pm 2.5\%$ for $|{\psi}\rangle_{A}=|{H}\rangle$ and $|{V}\rangle$, respectively. However, transferring coherent superpositions of $|H\rangle$ and $|V\rangle$ requires interference and thus exhibit more decoherence due to an imperfect spectro-temporal indistinguishability between Alice and Bob’s photons. For Alice transmitting $|{\psi}\rangle_{A}=|{D}\rangle$ and $|{A}\rangle$, we obtain fidelities of $F_{D}=87.5\pm 3.9\%$ and $F_{A}=85.5\pm 3.7\%$, respectively. Since the SpRS noise is unpolarized and independent of Alice’s prepared qubit, assuming symmetry of the remaining equatorial states we obtain an average fidelity for an arbitrary qubit on the Bloch sphere of $F_{\rm avg}=\frac{1}{3}F_{\text{poles}}+\frac{2}{3}F_{\text{equator }}=89.9\pm 3.1\%$. Furthermore, we measure $F_{\rm avg}=90.8\pm 0.8\%$ without sending the photons over the 30.2-km link (see supp. materials), demonstrating that there is minimal degradation of fidelity from distribution over the fiber nor the inclusion of the 400-Gbps signal. The state transfer fidelity of each qubit is shown in Fig. 4(e). All are well above the $F=2/3$ limit for classical-based methods [52], demonstrating that non-classical teleportation can be achieved alongside high-rate conventional communications.
@@ 4 Discussion and outlook
Our results also indicate the potential for integrating even more advanced quantum operations. For example, simple modifications to Alice’s photon pair source would generate entangled photon pairs and result in entanglement swapping [55], entangling initially independent photons at Alice and Bob’s nodes after the BSM. Based on this similarity, it is reasonable to conclude that swapping alongside high-rate classical data is achievable. Further, midpoint BSMs are important for applications such as networking with a central quantum processor [34], quantum repeaters [29], amongst others. Although these will require more capabilities beyond those demonstrated here (e.g., quantum memories, purification, etc.), our results are promising and could be useful for the ongoing development of this technology.
Although O-band teleportation networks show promise for high-fidelity integration, C-band quantum networking is usually preferred in quantum-dedicated fiber due to the lowest loss in silica fibers. Unfortunately, C-band channels suffer from orders of magnitude higher SpRS generation rates from C-band classical light [4, 17]. CV quantum encoding can help address this issue [25] but is much more fragile to loss compared to discrete-variables which limits its reach. One intriguing possibility is to use high-dimensional entanglement [56], but has yet to be fully investigated in high-power regimes. Using our theoretical model [49], we predict C-band teleportation fidelity would be significantly degraded with mW-level coexisting C-band power (see supp. materials) but would improve our fiber transmission loss by $\sim 0.15$ dB/km. This has relevance for research towards long-distance quantum backbone networks since conventional backbone fibers can have $\sim 100$ mW aggregate power. Further, WDM bandwidth may become limited as some conventional networks push towards fully populating the C- and L-bands [57], which also increases the difficulty of avoiding channel cross-talk and optical amplifier noise. One route to alleviate lower O-band quantum rates is to use wavelength multiplexing, which has been shown successful in allowing multi-channel entangled photons populating most of the O-band co-propagating with high-power C- and L-band WDM channels [17]. In any case, future research into novel methods to protect quantum fidelity from strong background noise would be invaluable.
@@
It is interesting to consider whether teleporting quantum information has any unique benefits in quantum-classical networking compared to direct transmission. It has been shown that using teleportation and BSM operations periodically along a long-distance fiber can increase the possible distances for quantum applications limited by detector dark count noise [28], however this may not necessarily apply for SpRS noise generated by classical light and depends on a vast number of parameters such as wavelength channel selections, fiber lengths, co- or counter-propagating signals, and optical amplification for long-distance classical communications. We predict that in some cases there is no advantage if the system has high SpRS noise, but using the methods applied here can help recover some improvements over long-distances and could in principle operate over 100’s of kms, albeit at impractically low quantum rates (see supp. materials). A full analysis of the various scenarios to consider will be pursued in future work. We will also investigate teleportation using a WCS source at Alice’s node [33, 36, 38, 39] as well as the consequences of multiplexing classical communications with entanglement swapping, quantum repeaters, and multi-photon entanglement.
In regards to real-world deployment, classical light-based signals are often used in quantum networking for synchronizing distant pulsed quantum sources and detectors to a common clock, communicating measurement results (e.g., BSM results), and monitoring environmental disturbances. Similarly, multiplexing these signals can enhance functionality without sacrificing potentially limited fiber resources [16, 59, 22, 21]. Since these can typically operate at sub-mW powers, our results show that including these in our system would have negligible impact if allocated to the C- or L-bands.